OpenStudy (anonymous):

What is the solution of n² – 49 = 0? –7 7 ±7 There is no solution.

OpenStudy (hoblos):

n² – 49 =0 (n-7)(n+7)=0 n=±7

hero (hero):

n² - 49 = 0 n² = 49 n = ± 7